# DepthVanish: Optimizing Adversarial Interval Structures for Stereo-Depth-Invisible Patches

### v1

@@ Abstract
Stereo Depth estimation is a critical task in autonomous driving and robotics, where inaccuracies (such as misidentifying nearby objects as distant) can lead to dangerous situations. Adversarial attacks against stereo depth estimation can help reveal vulnerabilities before deployment. Previous work has shown that repeating optimized textures can effectively mislead stereo depth estimation in digital settings. However, our research reveals that these naively repeated texture structures perform poorly in physical-world implementations, i.e., when deployed as patches, limiting their practical utility for testing stereo depth estimation systems. In this work, for the first time, we discover that introducing regular intervals between repeated textures, creating a striped structure, significantly enhances the patch attack effectiveness. Through extensive experimentation, we analyze how variations of this novel structure influence the performance. Based on these insights, we develop a novel stereo depth attack that jointly optimizes both the striped structure and texture elements. Our generated adversarial patches can be inserted into any scenes and successfully attack state-of-the-art stereo depth estimation methods, i.e., RAFTS-tereo and STTR. Most critically, our patch can also attack commercial RGB-D cameras (Intel RealSense) in real-world conditions, demonstrating their practical relevance for security assessment of stereo systems.
@@ 1 Introduction
Depth estimation is a critical component in safety-critical embodied systems like autonomous driving Cheng_2024_CVPR and robotics cai2024spatialbot , where accurate perception of the 3D environment is essential for reliable operation. Investigating the errors in depth estimation-such as mistaking nearby objects for distant ones in safety-critical embodied systems wong2020targeted ; chawla2022adversarial ; YunJCL23Poster ; ChengLCTCLZ22MDEAttack ; LiuWWHGX024Roadmarks ; ZhengHTZ0024piAttack ; ZhengLS00024cvpr -can provide critical insights for safety practice. Most existing work focuses on the security vulnerabilities of monocular depth estimation, which relies heavily on scene priors from single images. Stereo depth estimation, on the other hand, utilizes geometric constraints and typically provides more robust and metrically accurate results, making it attractive for high-stakes applications.
However, despite this inherent advantage, recent studies revealed that DNN-based stereo pipelines remain vulnerable to adversarial attacks, as carefully crafted pixel-level perturbations wong2021stereopagnosia ; berger2022stereoscopic can cause substantial disparity estimation errors. Nevertheless, previous research has primarily addressed digital attacks utilizing full-image noise, which are impractical in real-world contexts due to constraints such as limited patch size, varying viewing angles, and dynamic lighting conditions. As illustrated in the first row of Fig. 1, when applied as a physical patch, the Stereoscopic berger2022stereoscopic fails to effectively attack RAFT-Stereo and Intel RealSense. This lack of physically realizable and generalizable attack methods presents a significant limitation in evaluating the robustness of stereo systems, particularly as stereo depth estimation continues to be deployed in real-world, safety-critical applications.
@@
In this study, we address these limitations by introducing the first adversarial patch attack that is effective in both digital and physical settings against state-of-the-art (SOTA) deep stereo depth estimation models (Figure 1, second row). Fundamentally, we discover that adding regular intervals between repeated textures to form a spatial structure shows great potential for improving the attack effectiveness and enables digital-to-physical transferability. Through systematic analysis, we show how interval spacing and texture repetition influence attack success. These insights inform a novel optimization pipeline that jointly designs texture and structure to achieve high attack effectiveness across models and deployment settings. Thus, we propose a novel optimization pipeline that co-designs both texture elements and structural layout for generating adversarial patches that ❶ remain effective when physically printed and inserted into real scenes, ❷ work across diverse datasets and environments and ❸ generalize across different stereo depth estimation models, including commercial RGB-D sensors, i.e., Intel RealSense. In summary, our contributions are as follows,
  • We introduce the first adversarial attack that is both digitally and physically effective for deep stereo depth estimation models, including the SOTA RAFT-Stereo and Stereo Transformer.
  • Through empirical analysis, we discover that regular interval spacing between repeated textures significantly improves real-world transferability over naive texture repetitions.
  • We develop a joint optimization algorithm that co-designs the texture content and spatial structure of the patch to maximize the digital and physical attack effectiveness.
@@ 2 Related Work
Stereo depth estimation. Stereo-based depth estimation is a technique that infers scene depth from visual correspondences—captured as disparity maps—between pairs of stereo images in various applicable settings kittisceneflow ; yang2019drivingstereo ; VirtualKITTI2 ; Middlebury21 ; RamirezTPSMS22Booster . Traditional methods typically follow a multi-stage pipeline involving the computation of matching costs, cost aggregation, and optimization to predict and refine disparities ScharsteinS02 . In contrast, recent advances have incorporated deep neural networks Tosi2025 , enabling end-to-end learning of feature representations for correspondence matching and direct prediction of disparity and/or depth. In particular, CNN-based methods ChangC18PSMN ; Guo_2019_CVPR ; yang2019hierarchical ; ZhangPYT19GANet ; MaoLLDWKL21UASNet ; ShenDSRZZ22PCW-Net typically build 3D cost volumes from shared-weight feature encoders, while attention-based models LiLDDCTU21 ; GuoLYWTUYL22 ; SuJ22ChiTrans ; KaraevRGNV023DynamicStereo employ vision transformers to model global correspondences and disambiguate difficult regions. Compared to monocular methods depthanything2024 ; DepthAnythingV22024 , stereo offers improved robustness by leveraging geometric constraints from dual viewpoints, but still faces challenges in low-texture and repetitive areas zhao2024depthaware ; DBLP:journals/displays/LuSYD21 .
@@
Depth estimation attack. Thanks to their effectiveness and potential for real-time performance, depth estimation systems have become integral to safety-critical applications such as autonomous driving Cheng_2024_CVPR and robotic navigation cai2024spatialbot . Monocular depth estimation models, in particular, have been extensively studied under both digital wong2020targeted ; chawla2022adversarial ; YunJCL23Poster and physical adversarial attacks ChengLCTCLZ22MDEAttack ; LiuWWHGX024Roadmarks ; ZhengHTZ0024piAttack ; ZhengLS00024cvpr . These evaluations have revealed various system vulnerabilities and led to the development of tailored defense strategies hu2019analysis ; selfsupadvtraining2023/iclr ; selfsupadvtraining2024/tpami , including adversarial training and robust feature learning. In contrast, stereo depth estimation systems—despite their geometric soundness and widespread deployment—have received relatively limited adversarial attention. Existing research has focused primarily on digital, white-box attacks berger2022stereoscopic ; wong2021stereopagnosia , overlooking potential vulnerabilities in the physical world and leaving a lack of effective defense strategies. This gap is especially concerning given stereo systems’ reliance on precise left-right correspondence, where failures can have serious real-world implications—particularly in autonomous systems that depend on accurate and reliable 3D perception yang2019drivingstereo .
@@ 3.1 Naive Repetition Fails in Realistic Patch Attacks
Stereo depth estimation recovers 3D structure by identifying correspondences between left and right images, typically formulated as a pixel-wise optimization along epipolar lines:
@@
where $C(x,d)$ denotes the matching cost between pixel $x$ in the left image and pixel $x-d$ in the right image. When repetitive patterns are presented, the cost volume exhibits periodic ambiguity:
@@
Previous adversarial attacks inject repetitive noise over the entire image to exploit periodic ambiguities berger2022stereoscopic . Since global noise injection is impractical in real-world scenarios, we instead explore attacks using localized adversarial patches. Nevertheless, simply optimizing repetitive patterns within a patch does not yield effective attacks. As shown by the black curve in Fig. 2(a), simply repeating the pattern within patches results in predicted depth that remains same to the ground truth, indicating limited adversarial effectiveness. This is visually confirmed in Fig. 2(b) (top left), where a naive repeated patch yields a predicted depth of $7.1\,\mathrm{m}$, which is almost identical to the ground truth of $7\,\mathrm{m}$. This observation reveals a key limitation of existing studies: naive repetition fails to generate sufficient ambiguity within small patches, motivating the need for more structured pattern designs.
@@ 3.2 Structured Intervals: Enhancing Patch Adversarial Effectiveness
Figure 2: Adversarial effect of interval spacing on depth prediction. (a) Mean predicted depth (solid lines) and variance (shaded regions) for different interval strategies, averaged over interval widths of 2–10 px. The gray dashed band indicates $\pm$1.5 m from the ground truth. (b) Visualization of typical results where the physical ground truth depth is 7 m.
To address this limitation, we propose introducing structured intervals within repeated patterns to effectively amplify stereo matching ambiguity, thereby enhancing the adversarial effect of the patch.
We systematically evaluate the impact of different interval configurations on the RAFT-Stereo model using the KITTI dataset. As shown in Fig. 2(a) and visualized in Fig. 2(b), structured intervals notably enhance attack effectiveness: ❶ Basic Repeat (black): Depth predictions remain close to the ground truth, indicating minimal adversarial influence. ❷ Horizontal Interval (blue): Moderate overestimation beyond $15\,\mathrm{m}$ (i.e., a $20\,\mathrm{m}$ true depth yields a $\sim 24\,\mathrm{m}$ prediction). Visual results (Fig. 2(b), bottom left) confirm a slight increase to $7.2\,\mathrm{m}$. ❸ Vertical Interval (green): Produces larger errors, frequently reaching $\sim 30\,\mathrm{m}$ at a $20\,\mathrm{m}$ ground truth. In Fig. 2(b) (top right), the predicted depth surges to $12.7\,\mathrm{m}$. ❹ Grid Interval (red): Combining both directions produces the strongest adversarial effect, with depth predictions surpassing $40\,\mathrm{m}$ at a $23\,\mathrm{m}$ ground truth. In the visual result (Fig. 2(b), bottom right), the predicted depth reaches $14.1\,\mathrm{m}$, demonstrating a significant attack impact.
@@
Figure 3: RAFT-Stereo depth prediction performance under various interval structures and patch rotation degrees. (a) Illustration of rotation around the Y and X axes. (b) Depth prediction performance with different rotation degrees around X axis. (c) Depth prediction performance with different rotation degrees around Y axis.
@@ 3.3 Structured Intervals: Improving Attack Robustness across Viewpoints
A practical adversarial patch must maintain its effectiveness even when the patch is rotated or viewed from different orientations. This is particularly important under real-world deployment conditions, where precise placement is difficult to control. To this end, we systematically evaluate the impact of interval structure on attack robustness against patch rotation, as shown in Fig. 3.
For X-axis rotation (Fig. 3(b)): ❶ The horizontal (blue) and vertical (green) intervals exhibit angle-dependent performance, succeeding only at certain angles. ❷ The grid interval (red) is more robust, demonstrating more consistent effectiveness across different angles. For Y-axis rotation (Fig. 3(c)), although all configurations show moderate attack robustness across viewpoints, adding intervals still yields improvements. These results show that structured intervals improve attack robustness to patch rotation, which is essential for reliable adversarial attacks in real-world scenarios.
@@ 4 Problem Formulation
To formalize the stereo depth estimation task and define our attack objective, we begin with the following setup. Given a stereo image pair $(\mathbf{I}_{l},\mathbf{I}_{r})$ and $\mathbf{I}\in\mathbb{R}^{3\times H\times W}$ of a specific scene, a pretrained stereo depth estimation model $\mathcal{F}(\cdot)$ predicts the pixel-wise disparity map $\textbf{d}_{pred}=\mathcal{F}(\mathbf{I}_{l},\mathbf{I}_{r})\in\mathbb{R}^{H\times W}$. The corresponding depth map is computed as $\textbf{z}=\frac{f\times B}{\textbf{d}_{pred}}$ where $f$ and $B$ denote the focal length and baseline of the stereo camera rig respectively.
In general, the objective of the attack is to construct an adversarial patch $\mathbf{P}\in\mathbb{R}^{3\times h_{p}\times w_{p}}$ such that the stereo depth estimation model $\mathcal{F}(\cdot)$ produces an incorrect disparity output:
$\mathcal{F}(\mathbf{\hat{I}}_{l},\mathbf{\hat{I}}_{r})\neq\mathcal{F}(\mathbf{I}_{l},\mathbf{I}_{r})$ (3)
where $\mathbf{\hat{I}}_{l},\mathbf{\hat{I}}_{r}$ represent the stereo images with the adversarial patch $\mathbf{P}$ deployed at region $p$ in the original scene. As shown in Sec. 3, interval spacing can trigger critical estimation failures, such as the depth disappearance attack. To demonstrate this vulnerability, we define a stronger attack objective:
$\mathcal{F}(\mathbf{\hat{I}}_{l},\mathbf{\hat{I}}_{r})=\mathbf{0},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ s.t.\leavevmode\nobreak\ \textbf{d}^{p}_{gt}=\textbf{c},$ (4)
where the model predicts zero disparity for the patch region (i.e., infinite depth), despite the ground truth $\textbf{d}^{p}_{gt}$ indicating a fixed, close distance $\mathbf{c}$. This attack objective reveals more severe vulnerabilities than Eq. (3) and poses substantial safety risks, particularly when the patch is physically realizable and effective in real-world deployments.

### v2

@@ Abstract
Stereo depth estimation is a critical task in autonomous driving and robotics, where inaccuracies (such as misidentifying nearby objects as distant) can lead to dangerous situations. Adversarial attacks against stereo depth estimation can help reveal vulnerabilities before deployment. Previous works have shown that repeating optimized textures can effectively mislead stereo depth estimation in digital settings. However, our research reveals that these naively repeated textures perform poorly in physical implementations, i.e., when deployed as patches, limiting their practical utility for stress-testing stereo depth estimation systems. In this work, for the first time, we discover that introducing regular intervals among the repeated textures, creating a grid structure, significantly enhances the patch’s attack performance. Through extensive experimentation, we analyze how variations of this novel structure influence the adversarial effectiveness. Based on these insights, we develop a novel stereo depth attack that jointly optimizes both the interval structure and texture elements. Our generated adversarial patches can be inserted into any scenes and successfully attack advanced stereo depth estimation methods of different paradigms, i.e., RAFT-Stereo and STTR. Most critically, our patch can also attack commercial RGB-D cameras (Intel RealSense) in real-world conditions, demonstrating their practical relevance for security assessment of stereo systems. The code is officially released at: https://github.com/WiWiN42/DepthVanish
@@ 1 Introduction
Depth estimation is a crucial component in safety-critical embodied systems like autonomous driving Cheng_2024_CVPR and robotics cai2024spatialbot , where accurate perception of the 3D environment is essential for reliable operation. Investigating the errors in depth estimation, such as mistaking nearby objects as distant ones in safety-critical embodied systems wong2020targeted ; chawla2022adversarial ; YunJCL23Poster ; ChengLCTCLZ22MDEAttack ; LiuWWHGX024Roadmarks ; ZhengHTZ0024piAttack ; ZhengLS00024cvpr , can provide critical insights for safety practices. Most existing works focus on the security vulnerabilities of monocular depth estimation, which relies heavily on scene priors from single images. Stereo depth estimation, on the other hand, utilizes geometric constraints and typically provides more robust and metrically accurate results, making it attractive for high-stakes applications.
However, despite this inherent advantage, recent studies revealed that DNN-based stereo pipelines remain vulnerable to adversarial attacks, as carefully crafted pixel-level perturbations wong2021stereopagnosia ; berger2022stereoscopic can cause substantial disparity estimation errors. Nevertheless, previous works have primarily addressed digital attacks utilizing full-image noise, which are impractical in real-world contexts due to constraints like limited patch size, varying viewing angles, and dynamic lighting conditions, etc. As illustrated in the first row of Fig. 1, when applied as a physical patch, the existing Stereoscopic berger2022stereoscopic fails to effectively attack RAFT-Stereo and Intel RealSense. This lack of physically realizable and generalizable attack methods presents a significant limitation in evaluating the robustness of stereo systems, particularly as stereo estimation continues to be deployed in real-world, safety-critical applications.
@@
In this study, we address these limitations by introducing the first adversarial patch attack that is effective in both digital and physical settings against widely deployed deep stereo depth estimation models (Fig. 1 second row). Fundamentally, we discover that adding regular intervals among repeated textures to form a spatial structure shows great potential for improving the attack effectiveness and enables digital-to-physical transferability. Through systematic analysis, we show how interval spacing influences the attack success. These insights inform a novel optimization pipeline that jointly designs patches’ texture and structure to achieve high attack effectiveness across models and deployment settings. Thus, we propose a novel optimization pipeline that co-designs both texture elements and interval structure for generating adversarial patches that ❶ remain effective when physically printed and inserted into real scenes, ❷ work across diverse datasets and environments and ❸ generalize across different stereo depth estimation models, including commercial RGB-D sensors, i.e., Intel RealSense. In summary, our contributions are as follows,
  • We introduce the first adversarial attack that is both digitally and physically effective for deep stereo estimation models including the advanced RAFT-Stereo and Stereo Transformer.
  • By conducting a comprehensive empirical study, we discover that regular interval spacing among repeated textures significantly improves the patch attack effectiveness and its real-world transferability over naive texture repetitions.
  • We develop a joint optimization algorithm, i.e. DepthVanish, that co-designs the texture and its spatial structure within the patch to maximize the digital and physical attack effectiveness.
@@ 2 Related Work
Stereo depth estimation. Stereo-based depth estimation is a technique that infers scene depth from visual correspondences, which captured as disparity maps, between pairs of stereo images in various applicable settings kittisceneflow ; yang2019drivingstereo ; VirtualKITTI2 ; Middlebury21 ; RamirezTPSMS22Booster ; kong2024patch . Traditional methods typically follow a multi-stage pipeline involving the computation of matching costs, cost aggregation, and optimization to predict and refine disparities ScharsteinS02 ; ChangC18PSMN ; xu2020aanet ; duggal2019deeppruner . In contrast, recent advances have incorporated deep neural networks Tosi2025 , enabling end-to-end learning of feature representations for correspondence matching and direct prediction of disparity and/or depth. In particular, CNN-based methods ChangC18PSMN ; Guo_2019_CVPR ; yang2019hierarchical ; ZhangPYT19GANet ; MaoLLDWKL21UASNet ; ShenDSRZZ22PCW-Net typically build 3D cost volumes from shared-weight feature encoders, attention-based models LiLDDCTU21 ; GuoLYWTUYL22 ; SuJ22ChiTrans ; KaraevRGNV023DynamicStereo employ vision transformers to model global correspondences and disambiguate difficult regions, and iterative refinement methods lipson2021raft ; li2023crestereo ; xu2023igev apply a recurrent update operator to progressively converge on the final disparity, avoiding the memory-intensive 3D cost volumes. Compared to monocular methods depthanything2024 ; DepthAnythingV22024 , stereo offers improved robustness by leveraging geometric constraints from dual viewpoints, but still faces challenges in low-texture and repetitive areas zhao2024depthaware ; DBLP:journals/displays/LuSYD21 .
@@
Depth estimation attack. Due to their effectiveness and capability for real-time performance, depth estimation systems have become essential components of safety-critical applications such as autonomous driving Cheng_2024_CVPR and robotic navigation cai2024spatialbot . Monocular depth estimation models, in particular, have been extensively studied under both digital wong2020targeted ; chawla2022adversarial ; YunJCL23Poster and physical adversarial attacks ChengLCTCLZ22MDEAttack ; LiuWWHGX024Roadmarks ; ZhengHTZ0024piAttack ; ZhengLS00024cvpr . These evaluations have revealed various system vulnerabilities and led to the development of tailored defense strategies hu2019analysis ; selfsupadvtraining2023/iclr ; selfsupadvtraining2024/tpami , including adversarial training and robust feature learning. In contrast, despite their geometric soundness and widespread deployment, stereo depth estimation systems lipson2021raft ; li2021revisiting have received limited attention in adversarial research. Existing research has focused primarily on digital, white-box attacks berger2022stereoscopic ; wong2021stereopagnosia , overlooking potential vulnerabilities in the physical world. This gap is particularly concerning, as stereo systems rely on precise correspondence between left and right images. Failures in such systems can lead to serious consequences, especially in autonomous applications where accurate and reliable 3D perception yang2019drivingstereo is critical.
@@ 3.1 Naive Repetition Fails in Realistic Patch Attacks
Stereo depth estimation recovers 3D structure by identifying correspondences between left and right images zhou2022doublestar , typically formulated as a pixel-wise optimization along epipolar lines:
@@
where $d\in\mathbb{Z}$ represents the horizontal disparity between pixel $x$ in the left image and pixel xdx-d in the right image, and $C(x,d)$ denotes the matching cost between them. When repetitive patterns are presented, the cost volume exhibits periodic ambiguity ScharsteinS02 :
@@
Previous adversarial attacks berger2022stereoscopic ; wong2021stereopagnosia inject repetitive optimized noise over the entire image to exploit such periodic ambiguities. Since global injection is impractical in real-world scenarios, we instead explore attacks using localized adversarial patches. As shown by the black curve in Fig. 2(a), we deploy the repetitive noise from berger2022stereoscopic as patch into a real-world scene at different ground-truth depth and plot the corresponding predicted mean depth. It can be seen that simply repeating the noise within patches results in predicted depth that remains same to the ground truth, indicating limited adversarial effectiveness. This is visually confirmed in Fig. 2(b) (top left), where a naive repeated patch yields a predicted depth of $7.1\penalty 10000\ m$, which is almost identical to the ground truth of $7\penalty 10000\ m$. This observation reveals a key limitation of existing studies: naive repetition fails to generate sufficient ambiguity within practical patches, which motivates the need for more structured pattern designs.
@@ 3.2 Structured Intervals: Enhancing Patch Adversarial Effectiveness
Figure 2: Adversarial effect of interval spacing on depth prediction. (a) Mean predicted depth (solid lines) and variance (shaded regions) for different interval spacing strategies, averaged over interval widths of $2-10\penalty 10000\ px$. The gray dashed band indicates $\pm 1.5\penalty 10000\ m$ from the ground truth. (b) Visualization of depth prediction results for typical different interval spaced patches where the ground truth depth is $7\penalty 10000\ m$.
To address the limitation, we propose introducing regular intervals into the repetitive pattern to amplify the matching ambiguity as Eq.(2), thereby enhancing the adversarial effect of the patch.
As demonstrated in Fig. 2(b), given the patch with basic repetitive pattern from berger2022stereoscopic (top left), we add vertical (top right), horizontal (bottom left) and grid (bottom right) space to form patches with structured intervals. We systematically evaluate the impact of different interval configurations on the RAFT-Stereo model using the KITTI dataset. As shown in Fig. 2(a), structured intervals notably enhance attack effectiveness. ❶ Basic Repeat (black): the predicted patch depth remain close to the ground truth depth indicating minimal adversarial influence, which is also verified by the visualization in Fig. 2(b) (top left). ❷ Horizontal Interval (blue): moderate overestimation beyond $15\penalty 10000\ m$ (e.g., a $20\penalty 10000\ m$ true depth yields a $\sim 24\penalty 10000\ m$ prediction). Visual results (Fig. 2(b), bottom left) confirm a slight increase to $7.2\penalty 10000\ m$. ❸ Vertical Interval (green): produces larger errors, frequently reaching $\sim 30\penalty 10000\ m$ at a $20\penalty 10000\ m$ ground truth. In Fig. 2(b) (top right), the predicted depth surges to $12.7\penalty 10000\ m$. ❹ Grid Interval (red): combining intervals in both directions produces the strongest adversarial effect, with depth predictions surpassing $40\penalty 10000\ m$ at a $23\penalty 10000\ m$ ground truth. In the visual result (Fig. 2(b), bottom right), the predicted depth reaches $14.1\penalty 10000\ m$, demonstrating a significant adversarial effectiveness.
@@
Figure 3: RAFT-Stereo depth prediction performance under various interval structures and patch rotation degrees. (a) Illustration of rotation around the X and Y axes. (b) Depth prediction performance at different rotation degrees around X axis. (c) Depth prediction performance at different rotation degrees around Y axis.
@@ 3.3 Structured Intervals: Improving Attack Robustness across Viewpoints
A practical adversarial patch must maintain its effectiveness even when the patch is rotated or viewed from different orientations. This is particularly important under real-world deployment conditions, where precise placement is difficult to control. To this end, we systematically evaluate the impact of interval structure on attack robustness against patch rotation. As shown in Fig. 3(a), we rotate the patch along two axes (i.e., X and Y) and summarize the predicted mean depth in Fig. 3(b) and (c).
For X-axis rotation (Fig. 3(b)): ❶ the horizontal (blue) and vertical (green) intervals exhibit angle-dependent performance, succeeding only at certain angles; ❷ the grid interval (red) is more robust, demonstrating more consistent effectiveness across different angles. For Y-axis rotation (Fig. 3(c)), although all configurations show moderate attack robustness across viewpoints, adding intervals still yields improvements. These results show that structured intervals improve attack robustness to patch rotation, which is essential for reliable adversarial attacks in real-world scenarios.
@@ 4 Problem Formulation
To formalize the stereo depth estimation task and define our attack objective, we begin with the following setup. Given a stereo image pair $(\mathbf{I}_{l},\mathbf{I}_{r})$ where $\mathbf{I}_{l},\mathbf{I}_{r}\in\mathbb{R}^{3\times H\times W}$ of a specific scene, a pretrained stereo depth estimation model $\mathcal{F}(\cdot)$ predicts the pixel-wise disparity map $\textbf{d}_{pred}=\mathcal{F}(\mathbf{I}_{l},\mathbf{I}_{r})\in\mathbb{R}^{H\times W}$. The corresponding depth map is computed as $\textbf{z}=\frac{f\times B}{\textbf{d}_{pred}}$ where $f$ and $B$ denote the focal length and baseline of the stereo camera rig respectively.
In general, the objective of adversarial patch attack is to construct a patch $\mathbf{P}\in\mathbb{R}^{3\times h_{p}\times w_{p}}$ such that the stereo depth estimation model $\mathcal{F}(\cdot)$ produces an incorrect disparity output for the patch:
$\mathcal{F}^{p}(\mathbf{\hat{I}}_{l},\mathbf{\hat{I}}_{r})\neq\mathcal{F}^{p}(\mathbf{I}_{l},\mathbf{I}_{r})$ (3)
where $\hat{I}_{l}$ and $\hat{I}_{r}$ denote the stereo images with the adversarial patch $\mathbf{P}$, and $p$ indicates the corresponding pixel region occupied by the patch within the prediction results. As analyzed in Sec. 3, interval spacing can trigger critical depth estimation failures, i.e., the disappearance attack. To expose the severity of such vulnerability, we define a more destructive attack objective:
$\mathcal{F}^{p}(\mathbf{\hat{I}}_{l},\mathbf{\hat{I}}_{r})=\mathbf{0},\penalty 10000\ \penalty 10000\ \penalty 10000\ s.t.\penalty 10000\ \textbf{d}^{p}_{gt}=\textbf{c},$ (4)
where the model predicts zero disparity for the patch region (i.e., infinite depth), despite the ground truth disparity of the patch, $\textbf{d}^{p}_{gt}$, indicating a fixed, close distance $(f\times B)/\mathbf{c}$. This attack objective reveals more severe vulnerabilities than Eq. (3) and poses substantial safety risks, particularly when the patch is physically realizable and effective in real-world deployments.